### v1

# Affective Multimodal Agents with Proactive Knowledge Grounding for Emotionally Aligned Marketing Dialogue

Abstract
Recent advances in large language models (LLMs) have enabled fluent dialogue systems, but most remain reactive and struggle in emotionally rich, goal-oriented settings such as marketing conversations. To address this limitation, we propose AffectMind, a multimodal affective dialogue agent that performs proactive reasoning and dynamic knowledge grounding to sustain emotionally aligned and persuasive interactions. AffectMind combines three components: a Proactive Knowledge Grounding Network (PKGN) that continuously updates factual and affective context from text, vision, and prosody; an Emotion–Intent Alignment Model (EIAM) that jointly models user emotion and purchase intent to adapt persuasion strategies; and a Reinforced Discourse Loop (RDL) that optimizes emotional coherence and engagement via reinforcement signals from user responses. Experiments on two newly curated marketing dialogue datasets, MM-ConvMarket and AffectPromo, show that AffectMind outperforms strong LLM-based baselines in emotional consistency (+26%), persuasive success rate (+19%), and long-term user engagement (+23%), highlighting emotion-grounded proactivity as a key capability for commercial multimodal agents.
Figure 1: AffectMind: A proactive marketing dialogue system that senses user signals, aligns with emotion and intent, and generates empathetic, persuasive responses.
I Introduction
The rapid advancement of Large Language Models (LLMs) has revolutionised conversational AI, enabling dialogue systems to achieve an unprecedented level of linguistic fluency and contextual understanding [1, 2, 44]. However, despite these remarkable capabilities, most current dialogue agents remain fundamentally reactive, responding to user inputs without proactively guiding conversations toward specific goals [3]. This limitation becomes particularly pronounced in goal-oriented domains such as marketing, where successful interactions require not only understanding user intent but also recognizing emotional states, adapting persuasion strategies, and maintaining long-term engagement [4].
Traditional marketing dialogue systems face several critical challenges that limit their effectiveness in real-world applications. First, they often operate in a purely textual modality, missing crucial emotional cues from voice intonation, facial expressions, and body language that human sales representatives naturally leverage [5]. Second, these systems typically rely on static knowledge bases that cannot adapt to evolving product information, user preferences, or market trends in real-time [6]. Third, existing approaches lack sophisticated emotional intelligence, failing to recognize, understand, and appropriately respond to the complex emotional dynamics that characterize successful sales interactions [7].
The importance of emotional alignment in marketing conversations cannot be overstated. Research in consumer psychology has consistently shown that purchasing decisions are heavily influenced by emotional factors, with emotional connection to a brand or product often outweighing purely rational considerations [8, 9]. Moreover, customers expect increasingly personalized and empathetic interactions from AI systems, particularly in high-involvement purchase scenarios [10]. Current dialogue systems, however, struggle to maintain emotional coherence across extended conversations and often produce responses that feel mechanical or emotionally misaligned with user needs.
To address these fundamental limitations, we propose AffectMind, a novel multimodal affective dialogue agent designed specifically for emotionally intelligent marketing conversations. Our system introduces three key architectural innovations that work synergistically to achieve superior performance in commercial dialogue scenarios:
  • Proactive Knowledge Grounding Network (PKGN): A dynamic knowledge integration system that continuously updates both factual product information and emotional context representations by processing multimodal inputs including text, visual cues, and prosodic features. Unlike traditional static knowledge bases, PKGN enables real-time adaptation to new information and changing conversation dynamics.
  • Emotion-Intent Alignment Model (EIAM): A joint modeling framework that simultaneously processes user emotional states and purchase intentions, enabling dynamic adaptation of persuasion strategies. EIAM leverages advanced multimodal fusion techniques to create a unified representation of user affective and cognitive states.
  • Reinforced Discourse Loop (RDL): A reinforcement learning framework that optimizes long-term conversation outcomes by learning from user engagement signals, emotional responses, and ultimate conversion metrics. RDL enables the system to develop sophisticated conversation strategies that balance immediate emotional alignment with long-term persuasive goals.
Our main contributions can be summarized as follows:
  • In order to address the discrepancy between predominantly reactive language models and the requirement for emotionally aligned, goal-directed marketing dialogue, a comprehensive multimodal framework is proposed, namely AffectMind. This integrates Proactive Knowledge Grounding (PKGN), Emotion-Intent Alignment (EIAM), and a Reinforced Discourse Loop (RDL) to support sustained, persuasive interactions.
  • In order to address the paucity of realistic multimodal benchmarks for emotion-grounded marketing conversations, two large-scale datasets, MM-ConvMarket and AffectPromo, have been constructed, with rich annotations for emotional states, persuasion strategies, and conversion outcomes. The results demonstrate the significant superiority of AffectMind in comparison to strong LLM- and multimodal-based baselines across a range of evaluation metrics.
  • In order to address the practical and ethical challenges associated with the deployment of emotionally aware persuasive agents in commercial settings, a detailed analysis of transparency, privacy, fairness, and user agency considerations is provided. This analysis is then distilled into concrete deployment guidelines for responsible marketing dialogue systems.
The rest of this paper is organized as follows: Section II reviews related work in multimodal dialogue systems and affective computing. Section III presents our AffectMind framework in detail. Section IV describes our experimental setup and comprehensive evaluation results. Section V discusses implications, limitations, and future directions. Section VI concludes the paper with broader impact considerations.
II Related Work
II-A Multimodal Dialogue Systems
Recent advances in multimodal dialogue systems have focused on integrating visual and textual information to create more contextually aware conversational agents [11, 12]. Early multimodal dialogue work primarily combined pre-trained vision and language models through simple concatenation or attention mechanisms [13, 14]. More sophisticated approaches have emerged with the development of transformer-based architectures that can jointly process multiple modalities [15, 16].
Recent work has explored various fusion strategies for multimodal dialogue [17]. Early fusion approaches process raw multimodal inputs simultaneously [18], while late fusion methods combine modality-specific representations at higher levels [19]. Intermediate fusion techniques, such as cross-modal attention mechanisms [20], have shown promise in balancing computational efficiency with representation quality.
However, the majority of extant multimodal dialogue systems concentrate chiefly on visual question answering or general conversation scenarios, with limited consideration given to emotion-aware applications. Moreover, these systems generally function in a reactive manner, responding to user queries without proactively steering the conversation towards specific objectives. The present work addresses these limitations by introducing proactive reasoning capabilities specifically designed for emotionally rich marketing interactions.
II-B Affective Computing in Dialogue
Affective computing has emerged as a critical component of human-computer interaction, with particular relevance to dialogue systems [7, 21]. Early emotion recognition approaches focused on single modalities, such as text-based sentiment analysis [22] or speech emotion recognition [23]. More recent work has explored multimodal emotion recognition that combines facial expressions, voice characteristics, and linguistic content [24, 25].
In dialogue contexts, emotion recognition serves multiple purposes: understanding user emotional states, generating emotionally appropriate responses, and maintaining emotional coherence across conversations [26]. Several recent works have attempted to incorporate emotion awareness into dialogue generation. For example, MIME [27] uses emotion recognition to condition response generation, while EmpDialogue [28] focuses on empathetic response generation.
However, the majority of extant affective dialogue systems treat emotion as a secondary consideration rather than a primary driving force for conversation strategy. Prior studies also show that although empathetic or emotion-aware responses can improve user experience, they rarely integrate emotional factors into persuasion-oriented dialogue planning [46, 47]. Existing systems tend to prioritise emotional appropriateness over emotional persuasiveness, despite evidence that emotional expression and affective framing significantly influence persuasive effectiveness [48, 49, 50]. This oversight neglects the pivotal link between affective states and persuasive impact that typifies effective marketing and decision-making interactions.
II-C Knowledge Grounding and Proactive Dialogue
Knowledge grounding in dialogue systems involves integrating external knowledge sources to provide informative and factually accurate responses [6, 29]. Traditional approaches rely on static knowledge bases or retrieval systems that provide relevant facts based on conversation context [30, 31].
Recent work has explored dynamic knowledge integration approaches that can update knowledge representations during conversations [32, 33]. These systems typically use memory networks or attention mechanisms to selectively incorporate relevant knowledge while maintaining conversation coherence.
Proactive dialogue represents a significant departure from traditional reactive systems. Instead of merely responding to user inputs, proactive agents actively guide conversations toward specific objectives [34, 35]. This capability is particularly important in task-oriented domains such as customer service, sales, and education, where the agent must balance user needs with organizational goals.
However, existing proactive dialogue systems often lack sophisticated emotional modeling and struggle to maintain natural conversation flow while pursuing specific objectives. Our work addresses these limitations by introducing emotionally intelligent proactive strategies specifically designed for marketing contexts.
II-D Marketing Dialogue and Persuasive AI
The intersection of AI and marketing has generated significant research interest, particularly in the development of conversational marketing systems [10, 36]. Early marketing chatbots focused primarily on information provision and basic customer service functions [37]. More sophisticated systems have begun incorporating persuasion techniques derived from social psychology and behavioral economics [38, 39].
Persuasive AI systems face unique challenges related to user trust, transparency, and ethical considerations [40]. Recent work has explored various persuasion strategies in AI systems, including social proof, authority, scarcity, and reciprocity [41, 42]. However, most existing persuasive AI systems operate through simple rule-based approaches rather than sophisticated emotional intelligence.
The ethical implications of persuasive AI have received increasing attention from both researchers and policymakers [43, 45]. Key concerns include user manipulation, privacy protection, and the potential for AI systems to exploit vulnerable populations. Our work addresses these concerns through explicit ethical constraints and transparency mechanisms.
III Methodology
III-A Problem Formulation
Let $\mathcal{U}=\{u_{1},u_{2},...,u_{T}\}$ represent a sequence of user inputs in a marketing dialogue session, where each $u_{t}$ contains multimodal information including textual content $u_{t}^{text}$, visual features $u_{t}^{vision}$ (such as facial expressions and gestures), and prosodic features $u_{t}^{audio}$ (including tone, pitch, and speaking rate). The goal is to generate a sequence of system responses $\mathcal{R}=\{r_{1},r_{2},...,r_{T}\}$ that maximizes three key objectives:
$\max_{\mathcal{R}}\alpha\cdot\mathcal{E}(\mathcal{U},\mathcal{R})+\beta\cdot\mathcal{P}(\mathcal{U},\mathcal{R})+\gamma\cdot\mathcal{G}(\mathcal{U},\mathcal{R})$ (1)
where $\mathcal{E}(\mathcal{U},\mathcal{R})$ represents emotional alignment between user states and system responses, $\mathcal{P}(\mathcal{U},\mathcal{R})$ measures persuasive effectiveness toward marketing objectives, and $\mathcal{G}(\mathcal{U},\mathcal{R})$ quantifies long-term user engagement. The weights $\alpha$, $\beta$, and $\gamma$ balance these competing objectives while ensuring ethical constraints are maintained.
We formally define the emotional alignment objective as:
$\mathcal{E}(\mathcal{U},\mathcal{R})=\frac{1}{T}\sum_{t=1}^{T}\text{sim}(e_{t}^{user},e_{t}^{response})$ (2)
where $e_{t}^{user}$ and $e_{t}^{response}$ represent the emotional states of the user and the appropriateness of the system response at time $t$, respectively, and $\text{sim}(\cdot,\cdot)$ measures emotional compatibility.
Figure 2: Architecture of the AffectMind algorithm, showcasing multimodal fusion, proactive knowledge generation, emotion-intent alignment, and a reinforced discourse loop for adaptive marketing dialogues.
III-B AffectMind Architecture Overview
Figure 2 illustrates the overall architecture of AffectMind, which consists of three main components working in synergy to achieve emotionally aligned marketing dialogue. The system processes multimodal user inputs through parallel feature extraction pipelines, integrates this information through the PKGN knowledge grounding network, performs joint emotion-intent modeling via EIAM, and optimizes long-term conversation strategy through RDL.
The data flow begins with multimodal input processing, where textual content is encoded using a pre-trained language model (RoBERTa-large), visual features are extracted using a fine-tuned vision transformer, and audio features are processed using Wav2Vec2.0. These modality-specific representations are then fused through learned attention mechanisms and fed into the core AffectMind components.
III-C Proactive Knowledge Grounding Network (PKGN)
The PKGN module dynamically updates both factual and affective knowledge representations by processing multimodal inputs in real-time. Unlike traditional static knowledge bases, PKGN maintains two complementary knowledge representations: factual knowledge $K_{f}$ containing product information, specifications, and objective features; and affective knowledge $K_{a}$ capturing emotional associations, user preferences, and contextual mood indicators.
The knowledge update mechanism is formulated as:
$K_{t}=\text{Update}(K_{t-1},\text{Fuse}(f_{text},f_{vision},f_{audio}))$ (3)
where $K_{t}=[K_{f}^{t};K_{a}^{t}]$ represents the concatenated knowledge state at time $t$, and $\text{Fuse}(\cdot)$ combines multimodal features through a learned attention mechanism:
$\text{Fuse}(f_{text},f_{vision},f_{audio})=\text{MHA}([f_{text};f_{vision};f_{audio}])$ (4)
The multimodal fusion employs multi-head attention (MHA) to learn optimal combinations of modality-specific features. Each modality contributes differently to factual versus affective knowledge updates:
$\begin{split}K_{f}^{t}&=K_{f}^{t-1}+\alpha_{f}\cdot\text{FF}_{f}(\text{Fuse}(f_{text},f_{vision},f_{audio}))\\ K_{a}^{t}&=K_{a}^{t-1}+\alpha_{a}\cdot\text{FF}_{a}(\text{Fuse}(f_{text},f_{vision},f_{audio}))\end{split}$ (5)
where $\text{FF}_{f}$ and $\text{FF}_{a}$ are modality-specific feed-forward networks, and $\alpha_{f}$, $\alpha_{a}$ are learnable update rates.
The proactive knowledge selection mechanism identifies relevant knowledge based on predicted user needs and conversation trajectory:
$K_{selected}=\text{Attention}(Q_{context},K_{t},V_{t})$ (6)
where $Q_{context}$ represents the current conversation context, and the attention mechanism selects the most relevant knowledge for response generation.
III-D Emotion-Intent Alignment Model (EIAM)
EIAM jointly models user emotional states and purchase intentions to enable dynamic adaptation of persuasion strategies. The model architecture consists of two parallel encoding streams that process emotional and intentional signals, followed by a fusion network that creates unified user state representations.
The emotional encoding stream processes multimodal affective cues:
$E_{t}=\text{EmotionEncoder}(\text{Concat}(f_{facial},f_{prosodic},f_{linguistic}))$ (7)
where $f_{facial}$ represents facial expression features extracted from video input, $f_{prosodic}$ captures voice emotion indicators, and $f_{linguistic}$ encodes text-based sentiment and emotion markers.
The intent encoding stream focuses on purchase-related behavioral signals:
$I_{t}=\text{IntentEncoder}(\text{Concat}(f_{query},f_{behavior},f_{context}))$ (8)
where $f_{query}$ represents query intent features, $f_{behavior}$ captures user interaction patterns, and $f_{context}$ encodes conversation history and product context.
The emotion-intent fusion mechanism creates a unified representation that captures the relationship between affective states and purchase motivations:
$S_{t}=\text{FusionNetwork}([E_{t};I_{t};E_{t}\odot I_{t}])$ (9)
where $\odot$ represents element-wise multiplication to capture interaction effects between emotion and intent.
Dynamic strategy adaptation is achieved through a policy network that selects appropriate persuasion techniques based on the current user state:
$\pi_{t}=\text{softmax}(W_{\pi}S_{t}+b_{\pi})$ (10)
where $\pi_{t}$ represents the probability distribution over available persuasion strategies (e.g., logical appeal, emotional appeal, social proof, urgency creation).
III-E Reinforced Discourse Loop (RDL)
The RDL component implements a reinforcement learning framework that optimizes long-term conversation outcomes by learning from user engagement signals, emotional responses, and conversion metrics. The system treats each conversation turn as an action in a partially observable Markov decision process (POMDP), where the goal is to maximize expected cumulative reward.
The state representation at time $t$ combines user state, conversation context, and knowledge states:
$s_{t}=[S_{t};C_{t};K_{t}]$ (11)
where $S_{t}$ is the EIAM user state, $C_{t}$ represents conversation context features, and $K_{t}$ is the PKGN knowledge state.
The action space consists of discrete response strategies combined with continuous response parameters:
$a_{t}=[\text{strategy}_{t};\text{emotion\_tone}_{t};\text{information\_content}_{t}]$ (12)
The reward function incorporates multiple feedback signals with different time horizons:
$R_{t}=w_{1}R_{immediate}+w_{2}R_{engagement}+w_{3}R_{conversion}$ (13)
where $R_{immediate}$ provides immediate feedback based on user emotional response, $R_{engagement}$ measures sustained user interest, and $R_{conversion}$ represents ultimate marketing success.
The policy optimization follows a actor-critic approach with separate value and policy networks:
$\begin{split}\theta_{policy}&\leftarrow\theta_{policy}+\alpha\nabla_{\theta_{policy}}\log\pi_{\theta_{policy}}(a_{t}|s_{t})A_{t}\\ \theta_{value}&\leftarrow\theta_{value}+\beta\nabla_{\theta_{value}}(R_{t}-V_{\theta_{value}}(s_{t}))^{2}\end{split}$ (14)
where $A_{t}=R_{t}-V_{\theta_{value}}(s_{t})$ is the advantage function.
Algorithm 1 presents the complete training procedure for the Reinforced Discourse Loop component.
Algorithm 1 Reinforced Discourse Loop Training Algorithm
1:Initialize policy network $\pi_{\theta}$ and value network $V_{\phi}$
2:Initialize experience buffer $\mathcal{B}$
3:for each training episode do
4: Reset conversation state $s_{0}$
5:for $t=1$ to $T$ do
6:  Sample action $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$
7:  Execute action and observe user response
8:  Compute reward $R_{t}$ from user feedback
9:  Update state $s_{t+1}$
10:  Store $(s_{t},a_{t},R_{t},s_{t+1})$ in $\mathcal{B}$
11:end for
12: Compute advantage estimates $A_{t}$
13: Update policy using equation (12)
14: Update value function using equation (12)
15:end for
IV Experiments and Results
IV-A Experimental Setup
IV-A1 Implementation Details
Our implementation uses PyTorch 1.12 with CUDA 11.7 for GPU acceleration. Training was conducted on 4 NVIDIA A100 GPUs with 40GB memory each. The language model component uses RoBERTa-large (355M parameters) for text encoding, Vision Transformer (ViT-B/16) for visual feature extraction, and Wav2Vec2.0-large for audio processing. The total model contains approximately 1.2B parameters across all components.
Hyperparameter optimization was performed using Optuna with 100 trials. Key hyperparameters include: learning rate $lr=2\times 10^{-5}$ with cosine annealing, batch size of 16 conversations, maximum sequence length of 512 tokens, and dropout rate of 0.1. The reinforcement learning component uses discount factor $\gamma=0.95$ and GAE parameter $\lambda=0.95$.
IV-A2 Datasets
We evaluate AffectMind on two newly curated marketing dialogue datasets:
$M$-ConvMarket
: A comprehensive multimodal marketing conversation dataset containing 10,000 dialogue sessions with text, video, and audio recordings from real customer interactions across various product categories including electronics, fashion, home goods, and services. Each session contains an average of 15.3 turns and includes detailed annotations for emotional states, persuasion attempts, and conversion outcomes. The dataset was collected from volunteer participants interacting with human sales representatives, then post-processed to extract multimodal features and ground-truth labels.
AffectPromo
: A specialized dataset focusing on emotional dynamics in promotional conversations, featuring 5,000 annotated sessions with detailed emotion labels (using a 6-class emotion model: joy, sadness, anger, fear, surprise, disgust) and persuasion success indicators. This dataset emphasizes high-emotion scenarios such as limited-time offers, premium product sales, and customer retention conversations. Sessions average 22.1 turns and include rich contextual information about customer demographics, purchase history, and interaction preferences.
TABLE I: Dataset Statistics and Characteristics
Dataset Sessions Avg. Product Emotion Conversion
Turns Categories Classes Rate
$M$-ConvMarket 10,000 15.3 12 6 34.2%
AffectPromo 5,000 22.1 8 6 28.7%
Data preprocessing included conversation segmentation, multimodal feature alignment, and quality filtering to remove incomplete or corrupted sessions. Emotion annotations were validated through inter-annotator agreement studies achieving Cohen’s $(kappa>0.75)$. Conversion labels were determined based on actual purchase decisions within 7 days of conversations.
IV-A3 Baseline Methods
We compare AffectMind against several state-of-the-art baseline methods:
GPT-3.5 Baseline
: Fine-tuned GPT-3.5-turbo on marketing dialogue data with standard conversation prompting. This represents current industry practice for conversational AI in sales contexts.
GPT-4 Enhanced
: GPT-4 with carefully crafted prompts including emotion awareness instructions and sales methodology guidelines. This baseline tests whether sophisticated prompting can achieve comparable results to our specialized architecture.
MultiModal-BERT
: A BERT-based dialogue model enhanced with visual and audio feature fusion through concatenation and cross-modal attention mechanisms [16].
BLIP-2 Dialogue
: A recent vision-language model adapted for dialogue generation through fine-tuning on conversational data [12].
EmpDialogue++
: An enhanced version of the empathetic dialogue system with additional marketing-specific training and multimodal capabilities [28].
PersuaBot
: A specialized persuasive dialogue system implementing rule-based persuasion techniques from social psychology literature [4].
IV-A4 Evaluation Metrics
Our evaluation encompasses both objective computational metrics and subjective human assessments:
Emotional Consistency Score
: Measures alignment between predicted and actual user emotional responses using cosine similarity between emotion embeddings. Calculated as the average similarity across all conversation turns.
Persuasive Success Rate
: Percentage of conversations that result in successful conversions (purchases) within the evaluation period. This represents the primary business objective for marketing dialogue systems.
User Engagement Score
: Composite metric combining conversation length, user response rate, and interaction quality indicators such as follow-up questions and positive feedback signals.
Emotional Intelligence Quotient (EIQ)
: Novel metric measuring the system’s ability to recognize, understand, and appropriately respond to user emotions across different emotional states and transition scenarios.
Knowledge Integration Accuracy
: Measures how effectively the system incorporates relevant factual information while maintaining conversational flow and emotional appropriateness.
Response Quality
: Human-evaluated metric assessing naturalness, helpfulness, and appropriateness of system responses on a 5-point Likert scale.
IV-B Main Results
Figure 3: Performance Comparison on Marketing Dialogue Tasks
Performance Comparison on Marketing Dialogue Tasks
Table II presents the comprehensive performance comparison between AffectMind and baseline methods across all evaluation metrics. Our approach demonstrates significant improvements across all measured dimensions, with particularly strong performance in emotional consistency and user engagement. Results can be shown in Figure 3.
TABLE II: Performance Comparison on Marketing Dialogue Tasks
Method Emotional Persuasive User EIQ Response
Consistency Success (%) Engagement Score Quality
GPT-3.5 Baseline 72.3 64.1 58.7 3.2 3.1
GPT-4 Enhanced 76.8 67.3 62.4 3.6 3.5
MultiModal-BERT 75.8 67.3 61.2 3.4 3.3
BLIP-2 Dialogue 78.1 69.8 63.4 3.7 3.6
EmpDialogue++ 79.2 70.1 65.3 3.8 3.7
PersuaBot 74.6 71.2 59.8 3.3 3.4
AffectMind 91.1 76.3 72.1 4.3 4.2
Improvement +26% +19% +23% +13% +14%
Statistical significance testing using paired t-tests confirms that all improvements are statistically significant $p<0.001$. The emotional consistency improvement of 26% represents a substantial advancement in the system’s ability to maintain appropriate emotional tone throughout conversations. The persuasive success rate improvement of 19% translates to significant business value in real-world deployment scenarios.
Beyond the absolute numbers in Table II, we assess both statistical and practical significance. For statistical validity, we report means across multiple random seeds and conversation resamplings, and compute non-parametric bootstrap 95% confidence intervals; paired tests across identical conversation sets indicate consistent gains (Holm–Bonferroni corrected). For practical significance, two observations stand out. First, Emotional Consistency improvements translate into measurably smoother state transitions: the model is less likely to oscillate between contradictory tones within a 3–5 turn window, which directly stabilizes downstream persuasion strategies. Second, the Persuasive Success Rate gains are accompanied by higher user-initiated follow-ups, indicating that AffectMind’s improvements are not merely cosmetic.
We further examine calibration by comparing predicted emotion distributions to human labels via reliability curves. AffectMind exhibits lower Expected Calibration Error than baselines, suggesting better confidence–accuracy alignment. This matters in deployment since over-confident yet misaligned affect predictions tend to trigger inappropriate strategies.
A mediation analysis across sessions shows that improvements in EIQ partly mediate gains in User Engagement, which in turn mediate Persuasive Success. Qualitatively, AffectMind de-escalates negative states faster and sustains positive momentum longer; quantitatively, we observe that sessions with stable affect trajectories (low variance of turn-level affect deltas) are those with the highest conversion likelihood. This supports our design choice to optimize emotion–intent alignment before strategy selection.
TABLE III: Long Conversation Session Performance Stability Analysis
Dialogue Turns Emotional Consistency Knowledge Consistency Response Relevance User Engagement Memory Retention Strategy Effectiveness
1-10 91.5 93.2 94.1 88.7 95.3 89.6
11-20 90.8 91.7 92.8 86.4 91.8 87.3
21-30 89.3 89.5 90.2 83.1 87.4 84.9
31-40 87.6 86.8 87.9 79.8 82.7 81.2
41-50 85.2 83.4 84.6 75.3 77.9 77.8
51+ 82.7 79.1 80.3 70.6 72.5 73.4
Performance Degradation 9.6% 15.1% 14.7% 20.4% 23.9% 18.1%
IV-C Ablation Studies
We conducted comprehensive ablation studies to understand the contribution of each component and design choice. Table IV presents results for different component combinations, which was also depicted in Figure 4.
Figure 4: Ablation study results showing the impact of removing individual components (PKGN, EIAM, RDL) and design choices (multimodal input, dynamic knowledge) on emotional consistency, persuasive success, and user engagement.
TABLE IV: Ablation Study Results
System Configuration Emotional Persuasive User
Consistency Success (%) Engagement
AffectMind (Full) 91.1 76.3 72.1
- w/o PKGN 85.7 72.1 67.8
- w/o EIAM 83.2 69.4 65.2
- w/o RDL 88.3 71.9 68.5
- w/o Multimodal 79.4 68.7 61.3
- Static Knowledge 82.6 70.2 64.7
- Text Only 76.2 65.8 58.9
The ablation study reveals that each component contributes significantly to overall performance. The PKGN component shows the largest individual contribution to persuasive success, while EIAM most strongly impacts emotional consistency. The RDL component primarily improves user engagement through better long-term strategy optimization.
Multimodal input processing provides substantial benefits over text-only approaches, with visual emotion recognition contributing most significantly to emotional consistency improvements. Audio features, while less impactful individually, provide crucial disambiguation in emotionally ambiguous scenarios.
Table IV indicates that each module contributes materially. Removing the Proactive Knowledge Grounding Network (PKGN) most strongly hurts Persuasive Success, consistent with our hypothesis that timely and context-appropriate facts are the backbone of credible marketing dialogue. Removing the Emotion–Intent Alignment Module (EIAM) primarily degrades Emotional Consistency, confirming that accurate affect recognition and regulation are prerequisites for strategy selection. Disabling the Reinforcement Dialogue Learner (RDL) reduces User Engagement—the RL policy appears to learn when to pivot among strategies rather than merely which strategy to use.
We also examine interaction effects. PKGN+EIAM yields super-additive improvements relative to either alone: accurate affect tracking increases the marginal utility of freshly grounded knowledge, and the availability of precise, relevant facts reduces the cognitive dissonance users feel when tone and content diverge. These observations are consistent with the cross-attention fusion results in Table V, where richer inter-modal conditioning benefits both affect tracking and knowledge selection.
IV-D Qualitative Analysis
We qualitatively probe how AffectMind attains its quantitative gains by analyzing four axes: (i) multimodal fusion behavior, (ii) proactive knowledge grounding dynamics, (iii) emotion-intent alignment effects across affective states, and (iv) long-session stability.
TABLE V: Multimodal Fusion Strategy Performance Comparison
Performance Metrics Computational Metrics
Fusion Strategy Emot. Pers. User Infer. Params Fusion
Cons. Succ. Eng. Time (M) Eff.
Early Fusion 84.2 70.1 66.3 45 890 0.78
Late Fusion 82.7 68.9 64.8 38 875 0.82
Cross-Attention 91.1 76.3 72.1 52 920 0.75
Dynamic Gating 89.6 74.8 70.5 49 905 0.76
TABLE VI: PKGN Knowledge Update Mechanism Effectiveness Analysis
Knowledge Update Strategy Knowledge Relevance Response Accuracy Dialogue Coherence Information Timeliness User Satisfaction
Static Knowledge Base 0.72 0.68 0.75 0.61 3.2
Periodic Updates 0.81 0.74 0.79 0.73 3.6
Attention-based 0.85 0.79 0.82 0.78 3.9
PKGN (Dynamic) 0.93 0.87 0.91 0.89 4.3
TABLE VII: Emotion-Intent Alignment Effectiveness Analysis
Emotional State Intent Recognition Strategy Conversion Positive Emotional
Accuracy (%) Matching (%) Rate (%) Response Rate (%)
Positive 92.3 94.1 42.7 88.5
Neutral 87.6 85.3 31.2 76.8
Negative 83.4 79.8 18.9 65.3
Angry 78.1 72.6 12.4 58.7
Confused 85.9 83.2 25.7 71.4
Excited 90.8 91.5 38.9 85.2
Average 86.4 84.4 28.3 74.3
EIAM Enhanced 91.2 93.7 36.8 87.9
Figure 5: Performance Degradation Across Dialogue Turns
Table V shows a consistent pattern: cross-attention yields the strongest peak performance across Emot. Cons., Pers. Succ., and User Eng., while dynamic gating approaches this ceiling with lower average inference time. In human inspection, cross-attention produces turns that more tightly couple wording and nonverbal cues. Dynamic gating, in contrast, learns to defer multimodal reasoning when text and PKGN already provide high-confidence evidence, which explains its favorable efficiency without large quality loss.
Table III shows performance decay with length, dominated by memory retention (–23.9%) and engagement (–20.4%). Affect-aware compression that preserves (i) objections, (ii) explicit commitments, and (iii) affect-trend descriptors recovers a large share of the drop by keeping strategy selection grounded in what the user last cared about. Peaks align with objection and price-sensitivity turns; the model pivots from spec sheets to social proof (reviews) and logistics (return/warranty) before escalating to scarcity cues, mirroring the quantitative gains in Tables VI and VII. In a high-end electronics vignette, AffectMind de-escalates frustration by reframing “spec overload” into lifestyle outcomes, then grounds claims with verified review snippets; baselines persist with technical detail and lose the user.
V Discussion
V-A Implications for Marketing AI
The results demonstrate that emotion-grounded proactive dialogue represents a significant advancement in marketing AI capabilities. The 19% improvement in persuasive success rate could translate to substantial business impact in real-world deployments. For a typical e-commerce platform processing millions of customer interactions, this improvement could result in significant revenue increases.
The multimodal capabilities of AffectMind enable more natural and effective customer interactions compared to text-only systems. The ability to process visual and audio cues allows the system to detect emotional states that might not be explicitly expressed in text, leading to more appropriate and effective responses.
The proactive knowledge grounding capability addresses a critical limitation of current marketing chatbots. By dynamically updating knowledge representations based on conversation context and user behavior, the system can provide more relevant and timely information, leading to improved customer satisfaction and conversion rates.
V-B Ethical Considerations
The development of emotionally intelligent persuasive AI systems raises important ethical considerations that must be carefully addressed:
Transparency and User Awareness
: Users should be clearly informed when interacting with AI systems, particularly those designed to influence behavior. Our system includes explicit disclosure mechanisms and allows users to request information about the AI’s emotional and persuasive strategies.
Manipulation vs. Assistance
: There is a fine line between helpful persuasion and manipulative coercion. We implement explicit constraints to prevent the system from exploiting vulnerable emotional states or using deceptive practices. The system is designed to provide genuine value to users while achieving business objectives.
Privacy Protection
: The collection and processing of emotional data raises significant privacy concerns. Our system implements differential privacy mechanisms and provides users with control over their emotional data usage and storage.
Bias and Fairness
: Emotional AI systems may inadvertently discriminate against certain demographic groups or emotional expression styles. We conducted extensive bias testing and implemented fairness constraints to ensure equitable treatment across diverse user populations.
V-C Limitations and Future Work
Several limitations of the current work provide opportunities for future research:
Computational Efficiency
: The current implementation requires significant computational resources, limiting deployment scalability. Future work will explore more efficient architectures and optimization techniques for real-time deployment.
Cross-Cultural Generalization
: Emotional expression and persuasion effectiveness vary significantly across cultures. Current datasets primarily represent Western cultural contexts, and future work should explore cross-cultural adaptation mechanisms.
Long-Term Relationship Modeling
: Current evaluation focuses on individual conversation sessions. Future research should explore how emotionally intelligent systems can build and maintain long-term customer relationships across multiple interactions.
Multi-Party Conversations
: Real-world marketing scenarios often involve multiple participants. Future work should extend the framework to handle complex multi-party dynamics and group decision-making processes.
Adversarial Robustness
: The system’s vulnerability to adversarial attacks and manipulation attempts requires further investigation. Future research should develop robust defense mechanisms against potential misuse.
VI Conclusion and Future Work
In this paper, we presented AffectMind, a novel multimodal affective dialogue agent designed for emotionally aligned marketing conversations. Our approach introduces three key architectural innovations: the Proactive Knowledge Grounding Network (PKGN) for dynamic knowledge integration, the Emotion-Intent Alignment Model (EIAM) for joint emotional and intentional reasoning, and the Reinforced Discourse Loop (RDL) for sustained engagement optimization.
The comprehensive experimental evaluation demonstrates significant improvements over state-of-the-art baselines across multiple metrics: emotional consistency (+26%), persuasive success rate (+19%), and user engagement (+23%). These results establish emotion-grounded proactivity as a crucial capability for next-generation conversational AI systems in commercial applications.
Our contributions extend beyond technical improvements to include important considerations for ethical AI deployment in persuasive contexts. The explicit attention to transparency, user agency, and fairness provides a framework for responsible development of emotionally intelligent commercial AI systems.
The newly introduced datasets (MM-ConvMarket and AffectPromo) provide valuable resources for future research in multimodal marketing dialogue. These datasets capture the complexity of real-world commercial interactions and enable standardized evaluation of emotional AI capabilities.
Future research directions include extending AffectMind to other goal-oriented dialogue domains, improving computational efficiency for large-scale deployment, developing more sophisticated ethical safeguards, and exploring cross-cultural adaptation mechanisms. Additionally, investigation of long-term relationship building and multi-party conversation dynamics represents promising areas for advancement.
The broader impact of this work extends to the transformation of customer-business interactions through more natural, empathetic, and effective AI-mediated conversations. As AI systems become increasingly integrated into commercial contexts, the principles and techniques developed in this work will be crucial for creating beneficial outcomes for both businesses and customers.